# V-RoAst: Visual Road Assessment Can VLM be a Road Safety Assessor Using the iRAP Standard?

### v4

@@
[…]
V-RoAst: Visual Road Assessment
Can VLM be a Road Safety Assessor Using the iRAP Standard?
Natchapon Jongwiriyanurak1,4, Zichao Zeng1,4,5, June Moh Goo1,5, Xinglei Wang1,4, Ilya Ilyankou1,4,
Kerkritt Sriroongvikrai2, Nicola Christie1, Meihui Wang1,4, Huanfa Chen3, James Haworth1,4
1 Department of Civil, Environmental and Geomatic Engineering (CEGE), UCL
2 Department of Civil Engineering, Chulalongkorn University
3 Centre for Advanced Spatial Analysis (CASA), UCL 4 SpaceTimeLab 5 3DIMPACT
Corresponding author: natchapon.jongwiriyanurak.20@ucl.ac.ukCorresponding author: j.haworth@ucl.ac.uk
Road safety assessments are critical yet costly, especially in Low- and Middle-Income Countries (LMICs), where most roads remain unrated. Traditional methods require expert annotation and training data, while supervised learning-based approaches struggle to generalise across regions. In this paper, we introduce V-RoAst, a zero-shot Visual Question Answering (VQA) framework using Vision-Language Models (VLMs) to classify road safety attributes defined by the iRAP standard. We introduce the first open-source dataset from ThaiRAP, consisting of over 2,000 curated street-level images from Thailand annotated for this task. We evaluate Gemini-1.5-flash and GPT-4o-mini on this dataset and benchmark their performance against VGGNet and ResNet baselines. While VLMs underperform on spatial awareness, they generalise well to unseen classes and offer flexible prompt-based reasoning without retraining. Our results show that VLMs can serve as automatic road assessment tools when integrated with complementary data. This work is the first to explore VLMs for zero-shot infrastructure risk assessment and opens new directions for automatic, low-cost road safety mapping. Code and dataset: https://github.com/PongNJ/V-RoAst.
@@ 1 Introduction
To reduce the cost of manual road assessments, automated detection of road features from imagery has become a widely adopted approach—typically using Convolutional Neural Networks (CNNs). While more affordable than human labelling, these models require task-specific training data, making them time-consuming to scale across regions due to visual variability in road environments across cities and countries [23, 34, 21, 38, 37]. Other studies have explored alternative data sources to overcome these limitations, including LiDAR [5], satellite imagery [6, 1], UAV imagery [4], and GPS traces [46].
@@ 2 Literature review
VLMs have been applied to various spatial tasks, including building understanding and captioning [11, 27], and autonomous driving [42, 20, 49, 45]. While their use in urban analytics is growing, applications in policy-driven, structured classification tasks—such as iRAP-based road assessments—remain largely unexplored.
@@ 4.2 Visual Road Assessment Framework
V-RoAst is a framework designed to classify 52 iRAP attributes $\mathcal{A}$ from the input image $I$, as shown in Figure 4. This framework is easily applicable in any city, requiring minimal expertise. The framework utilises VLMs via text input for system instructions $T_{S}$ and user prompts $T_{U}$ which include $I$ and $M$. The workflow operates by inputting an image $I$ and its associated metadata $M$, including task specifications, local context, attribute details, and output format into the VLM $f$. This process enables the generation of attribute-level predictions $\mathcal{A}$ for each image $I$.
@@ 4.2.1 System Instruction
    Task Specification: This component provides step-by-step instructions for processing each image. The model is guided to classify each <<<Attribute>>> by selecting the most appropriate <<<Category class>>> based on its visual description. This section also specifies how the output should be formatted. Notably, {country} is included to localise the analysis for each region.
@@
    Local Context: Information such as which side of the road vehicles drive on is provided through {local_context}. This helps the VLMs interpret images more accurately by incorporating local traffic rules and infrastructure norms. It is especially useful for visually ambiguous attributes—like the presence of service roads—where local knowledge improves classification. Including this input mimics how human coders use familiarity with the area to make more accurate assessments.
@@
    Attribute Details: Each attribute is defined with its name, description ($<$Attribute description$>$), and possible classes ($<$Category class$>$), accompanied by descriptions of each category ($<$Category description$>$). These structured definitions ensure that the models have clear guidelines for classification. Note that this section includes approximately 100,000 characters (or 20,000 tokens) from the iRAP coding manual, which this work made available in JSON format.
@@
    Output Format: The results are returned JSON format to maintain consistency. The models are instructed to select the best match from the provided $<$Category class$>$ list, making the outputs standardised and interpretable.
@@ 4.2.2 User Prompt
Speed-related attributes such as “Speed Limit,” “Motorcycle Speed Limit,” and “Speed Management” are essential for iRAP assessments but are often not directly observable from imagery alone. To address this, our framework incorporates geographic metadata (e.g., GPS coordinates) into the VLM prompt, enabling the model to associate locations with region-specific traffic regulations learned from training data. For instance, rural roads in certain provinces may visually resemble one another but differ in legal speed limits— context, the model learns to infer through spatial priors. By combining visual input with location-aware prompts, VLMs draw on both scene appearance and broader regulatory patterns seen during pretraining. This multimodal design improves the model’s ability to classify speed-related attributes with greater accuracy and contextual relevance, especially when explicit signage is absent.
@@ 4.4 Image Processor for Mapillary Imagery
Crowdsourced Street View Images (SVIs) are accessible on various platforms, with Mapillary being one of the most well-known, providing an API for image downloads. Hence, we evaluate the system using the V-RoAst framework with Gemini-1.5-flash to classify all 52 attributes. Then, we combine these results with additional information—including operating speed and ADDT from the ground truth—to determine the star rating outcomes.
@@
It is important to note that some attributes may differ from the iRAP ground truth. For example, the number of vehicles shown in the image (observed flow) and the number of vehicles parked in the captured scene can vary. To validate the automatic use of V-RoAst, we did not verify whether the downloaded images originated from the exact same location, which could potentially affect the predicted star rating.
@@ 5.1 Implementation Detail
We report results at two granularities. Overall performance is the macro-average of each metric across the iRAP attribute groups (“sections”), summarising model behaviour over broad feature types. Attribute-level performance is computed for each of the 52 individual attributes, revealing fine-grained strengths and weaknesses. When a road segment carries multiple risk classes, we follow iRAP convention and keep the first (highest-risk) label. For every attribute we calculate accuracy, precision, recall, and F1-score, then macro-average them so that attributes with few samples contribute equally to the final score.
@@ 5.2.1 Overall Performance
In contrast, the zero-shot VLMs show reduced performance. Gemini-1.5-flash achieves a macro-accuracy of 0.82 and an F1-score of 0.47, while GPT-4o-mini performs slightly behind. Despite this performance gap, VLMs offer a crucial advantage—the ability to predict all possible iRAP attribute classes without requiring task-specific training data. This is particularly valuable in data-sparse or label-scarce settings.
@@ 5.2.3 Attribute-Level Analysis
Table 3 in the Appendix illustrates per-attribute performance. Baseline models were trained only on attributes with sufficient class diversity, excluding 11 attributes with a single class. VLMs are not constrained by this limitation and generate predictions for all attributes, offering broader applicability.
We observe that VLMs perform better when recognising prominent visual cues (e.g., presence of a school zone, number of lanes) but underperform in estimating distances or interpreting scene geometry. These results suggest that while VLMs are effective for attribute presence/absence tasks, they may struggle with precise spatial reasoning [36].
@@ 5.2.4 Qualitative Assessment using VQA
A key advantage of VLMs lies in their ability to perform VQA, enabling users–regardless of technical background–to iteratively refine model outputs through prompt adaptation. This makes VLMs particularly suitable for participatory or practitioner-led applications, where model interpretability and adaptability are crucial.
Figure 6 illustrates how prompt tuning can be employed to guide model reasoning and address misclassification. For example, in the bottom-right case, none of the four models tested correctly identified the “Sidewalk (passenger-side)” attribute. The ground truth label referred to an informal path located more than 1 metre from the main carriageway, a feature not clearly visible in the street-view image. This discrepancy highlights a common challenge that visual cues may be insufficient or ambiguous, requiring local contextual knowledge for accurate classification.
@@ 6 Conclusion and Future work
This work introduced V-RoAst, a zero-shot prompting approach VLMs such as Gemini-1.5-flash and GPT-4o-mini for road safety assessment. Since VLMs enable flexible prompt engineering and results in potential adaptability improvements which is crucial for resource-limited settings in LMICs. VLMs excel at simple, interpretable attributes but struggle with spatially complex or implicit features like informal sidewalks. Limited street view coverage also impacts accuracy, highlighting the potential benefit of integrating satellite imagery for added context.
@@
Appendix
@@
Observed Flow Motorcycle (OF) 0.96 0.96 0.80 0.93 0.48 0.67 0.20 0.58 0.48 0.97 0.27 0.42 0.48 0.73 0.20 0.46
@@
Speed Limits Speed limit 0.99 0.98 0.44 0.74 0.65 0.98 0.34 0.50 0.67 0.97 0.33 0.48 0.66 0.98 0.30 0.48
@@
Mid-block Number of lanes 0.95 0.95 0.14 0.55 0.66 0.82 0.29 0.30 0.44 0.82 0.32 0.54 0.49 0.82 0.09 0.26

### v5

@@
Abstract
Road safety assessments are critical yet costly, especially in Low- and Middle-Income Countries (LMICs), where most roads remain unrated. Traditional methods require expert annotation and training data, while supervised learning-based approaches struggle to generalise across regions. In this paper, we introduce V-RoAst, a zero-shot Visual Question Answering (VQA) framework using Vision-Language Models (VLMs) to classify road safety attributes defined by the iRAP standard. We introduce the first open-source dataset from ThaiRAP, comprising over 2,000 curated street-level images from Thailand, annotated for this task. We evaluate Gemini-1.5-flash and GPT-4o-mini on this dataset and benchmark their performance against VGGNet and ResNet baselines. While VLMs underperform on spatial awareness, they generalise well to unseen classes and offer flexible prompt-based reasoning without retraining. Our results show that VLMs can serve as automatic road assessment tools when integrated with complementary data. This work is the first to explore VLMs for zero-shot infrastructure risk assessment and opens new directions for automatic, low-cost road safety mapping. Code and dataset: https://github.com/PongNJ/V-RoAst.
@@ 1 Introduction
To reduce the cost of manual road assessments, automated detection of road features from imagery has become a widely adopted approach, typically using Convolutional Neural Networks (CNNs). While more affordable than human labelling, these models require task-specific training data, making them time-consuming to scale across regions due to visual variability in road environments across cities and countries [23, 34, 21, 38, 37]. Other studies have explored alternative data sources to overcome these limitations, including LiDAR [5], satellite imagery [6, 1], UAV imagery [4], and GPS traces [46].
@@ 2 Literature review
VLMs have been applied to various spatial tasks, including building understanding and captioning [11, 27], and autonomous driving [42, 20, 49, 45]. While their use in urban analytics is growing, applications in policy-driven, structured classification tasks, such as iRAP-based road assessments, remain largely unexplored.
@@ 4.2 Visual Road Assessment Framework
V-RoAst is a framework designed to classify 52 iRAP attributes $\mathcal{A}$ from the input image $I$, as shown in Figure 4. This framework is easily applicable in any city, requiring minimal expertise. The framework uses VLMs via text input for system instructions $T_{S}$ and user prompts $T_{U}$, which include $I$ and $M$. The workflow operates by inputting an image $I$ and its associated metadata $M$, including task specifications, local context, attribute details, and output format into the VLM $f$. This process enables the generation of attribute-level predictions $\mathcal{A}$ for each image $I$.
@@ 4.2.1 System Instruction
    Task Specification: This component provides step-by-step instructions for processing each image. The model is guided to classify each <<Attribute>> by selecting the most appropriate <<Category class>> based on its visual description. This section also specifies how the output should be formatted. Notably, {country} is included to localise the analysis for each region.
@@
    Local Context: Information such as which side of the road vehicles drive on is provided through {local_context}. This helps the VLMs interpret images more accurately by incorporating local traffic rules and infrastructure norms. It is especially useful for visually ambiguous attributes, like the presence of service roads, where local knowledge improves classification. Including this input mimics how human coders use familiarity with the area to make more accurate assessments.
@@
    Attribute Details: Each attribute is defined with its name, description (<<Attribute description>>), and possible classes (<<Category class>>), accompanied by descriptions of each category (<<Category description>>). These structured definitions ensure that the models have clear guidelines for classification. Note that this section includes approximately 100,000 characters (or 20,000 tokens) from the iRAP coding manual, which this work made available in JSON format.
@@
    Output Format: The results are returned in JSON format to maintain consistency. The models are instructed to select the best match from the provided <<Category class>> list, making the outputs standardised and interpretable.
@@ 4.2.2 User Prompt
Speed-related attributes such as “Speed Limit” and “Motorcycle Speed Limit” are essential for iRAP assessments but are often not directly observable from imagery alone, particularly when speed signage is absent or unclear. In V-RoAst, these attributes are predicted by providing the VLM with both the image and its coordinates in the prompt. The location information enables the model to draw on spatial priors learned during pretraining (e.g., typical speed regulations in that region) while also inferring additional contextual cues, such as road type, land use, and traffic environment, from the image itself. By combining these two sources, the model can estimate the most likely speed limit even when explicit signage is missing. For example, two visually similar rural roads may have different statutory limits depending on the province; the geographic cue guides the model toward the correct classification. This approach improves the contextual relevance and accuracy of speed-related predictions without requiring explicit annotation of road type or land use.
@@ 4.4 Image Processor for Mapillary Imagery
Crowdsourced Street View Images (SVIs) are accessible on various platforms, with Mapillary being one of the most well-known, providing an API for image downloads. Hence, we evaluate the system using the V-RoAst framework with Gemini-1.5-flash to classify all 52 attributes. Then, we combine these results with additional information, including operating speed and ADDT from the ground truth, to determine the star rating outcomes.
@@
It is important to note that some attributes may differ from the iRAP ground truth. For example, the number of vehicles shown in image (observed flow) and the number of vehicles parked in the captured scene can vary. To validate the automatic use of V-RoAst, we did not verify whether the downloaded images originated from the same road, which could potentially affect the predicted star rating.
@@ 5.1 Implementation Detail
We report results at two granularities. Overall performance is the macro-average of each metric across the iRAP attribute groups (“sections”), summarising model behaviour over broad feature types. Attribute-level performance is computed for each of the 52 individual attributes, revealing fine-grained strengths and weaknesses. When a road segment carries multiple risk classes, we follow the iRAP convention and keep the first (highest-risk) label. For every attribute, we calculate accuracy, precision, recall, and F1-score, then macro-average them so that attributes with few samples contribute equally to the final score.
@@ 5.2.1 Overall Performance
In contrast, the zero-shot VLMs show reduced performance. Gemini-1.5-flash achieves a macro-accuracy of 0.82 and an F1-score of 0.47, while GPT-4o-mini performs slightly behind. Despite this performance gap, VLMs offer a crucial advantage, the ability to predict all possible iRAP attribute classes without requiring task-specific training data. This is particularly valuable in data-sparse or label-scarce settings.
@@ 5.2.3 Attribute-Level Analysis
Attribute-level performance analysis shows that baseline models were trained only on attributes with sufficient class diversity, excluding 11 attributes that contained a single class. VLMs are not constrained by this limitation and can generate predictions for all attributes, offering broader applicability. Full attribute-level results are provided in the Supplementary Material.
We observe that VLMs perform better when recognising prominent visual cues (e.g., vehicle parking, area type) but underperform in estimating distances or interpreting scene geometry. These results suggest that while VLMs are effective for attribute presence/absence tasks, they may struggle with precise spatial reasoning [36].
@@ 5.2.4 Qualitative Assessment using VQA
A key advantage of VLMs lies in their ability to perform VQA, enabling users, regardless of technical background, to iteratively refine model outputs through prompt adaptation. This makes VLMs particularly suitable for participatory or practitioner-led applications, where model interpretability and adaptability are crucial.
Figure 6 illustrates how prompt tuning can be employed to guide model reasoning and address misclassification. For example, in the bottom-right case, none of the four models tested correctly identified the “Sidewalk - passenger-side” attribute. The ground truth label referred to an informal path located more than 1 metre from the main carriageway, a feature not clearly visible in the street-view image. This discrepancy highlights a common challenge that visual cues may be insufficient or ambiguous, requiring local contextual knowledge for accurate classification.
@@ 6 Conclusion and Future work
This work introduced V-RoAst, a zero-shot prompting approach to VLMs such as Gemini-1.5-flash and GPT-4o-mini for road safety assessment. Since VLMs enable flexible prompt engineering and result in potential adaptability improvements, which is crucial for resource-limited settings in LMICs. VLMs excel at simple, interpretable attributes but struggle with spatially complex or implicit features like informal sidewalks. Limited street view coverage also impacts accuracy, highlighting the potential benefit of integrating satellite imagery for added context.
@@
  • Geometry reasoning: Investigate techniques like depth estimation or segmentation to handle attributes requiring spatial measurements.
@@
Acknowledgements
We would like to thank the Thai Road Assessment Programme (ThaiRAP) and Chulalongkorn University for providing the dataset used in this work. Support for this work was provided in part by the Regan Smeed Travel Fund from the UCL Centre for Transport Studies (CTS). Natchapon Jongwiriyanurak is funded by Royal Thai Government Scholarship. Zichao Zeng, June Moh Goo and Ilya Ilyankou are funded by Ordnance Survey and UKRI Engineering and Physical Sciences Research Council (EPSRC) [Grant numbers EP/W522077/1, EP/X524840/1 and EP/Y528651/1]. Xinglei Wang is jointly funded by UCL Dean’s Prize and China Scholarship Council [Grant number 202106270039].
Supplemental Material
@@ Acknowledgements
Observed Flow
Motorcycle (OF) 0.96 0.96 0.80 0.93 0.48 0.67 0.20 0.58 0.48 0.97 0.27 0.42 0.48 0.73 0.20 0.46
@@
Speed Limits
Speed limit 0.99 0.98 0.44 0.74 0.65 0.98 0.34 0.50 0.67 0.97 0.33 0.48 0.66 0.98 0.30 0.48
@@
Mid-block
Number of lanes 0.95 0.95 0.14 0.55 0.66 0.82 0.29 0.30 0.44 0.82 0.32 0.54 0.49 0.82 0.09 0.26
@@
This supplementary section expands on the attribute-level analysis introduced in Section 5.2.3 of the main paper. Table 3 provides detailed performance results for all 52 iRAP-defined attributes, comparing four models, two CNN-based baselines (VGG and ResNet) and two vision–language models (GPT-4o-mini and Gemini-1.5-flash).
Metrics reported include Accuracy, Precision, Recall, and F1-score. The best-performing model for each metric is bolded. This breakdown complements the summary metrics in the main paper and highlights performance variation across diverse attribute types.
As discussed, vision–language models are capable of zero-shot prediction across all attributes, including those with limited or no training samples in the CNN setting.